### 21 Feb 2024 v1

# Bound states in the continuum in a tangential ring with pointlike impurities

Abstract
Quantum rings coupled to external nanowires offer a versatile platform for the manipulation of the quantum mesoscopic transport. We here study such a system including periodically distributed pointlike impurities along the ring. Based on an exact expression for the conductance found here, we demonstrate that the bound-states in continuum (BICs) form from the ring states at the high-symmetry momenta in the ring’s Brillouin zone. Furthermore, the presence of the inversion symmetry allows for a selective decoupling of resonant states, favoring the BIC generation and therefore allowing an extra tunability in the quantum transport of the system. Finally, we suggest that the magnetic fluxes and Rashba spin-orbit coupling offer other possible routes for the BIC formation in laterally coupled quantum rings.
Introduction. Bound states in the continuum (BICs) remain localized and coexist with a continuous spectrum of radiating states that propagate outside the system [1, 2]. Although this special class of states was proposed at the beginning of quantum mechanics by von Neumann and Wigner [3], only in the following decades such states became the subject of more intense research [4, 5]. Furthermore, the concept of BICs is not exclusively operative in quantum mechanics but also pertains to classical wave phenomena, as they have been shown to occur in various setups, e.g., in photonics, acoustics, and electronics [6, 7, 8, 9, 10]. Due to the advancement in nanofabrication technology and their experimental observation [11, 12], BICs have recently attracted considerable interest for applications, as in building lasers, sensors, and filters [13, 14, 15, 16, 17, 18].
In mesoscopic systems, BICs can be formed by several mechanisms, with destructive interference in resonant states possibly being the most prominent one [1]. In this respect, quantum rings play a rather special role due to their geometry, also offering flexibility of manipulation, as they have already been intensively studied both theoretically [19, 20, 21, 22, 23, 24, 25] and experimentally [26, 27, 28, 29, 30, 31, 32, 33]. Moreover, mesoscopic coupled waveguide systems feature various sources of disorder, often modeled as point-like impurities or short-range scatterers [19, 20, 21, 22]. On the other hand, in such coupled mesoscopic systems, the appearance of a subset of discrete energy levels mixed with a continuum of states is attributed to the Fano effect [34], manifesting as an asymmetric peak-and-dip feature in the conductance. However, the case of many impurities has not been intensively studied despite the expected richness in the transport behavior. In particular, a strong dependence of the conductance on the distribution of impurities and the system’s geometry is expected primarily due to quantum interference phenomena dominating the physics at the mesoscopic scales. Furthermore, the theory of electronic transport, relating the conductance to the transmission probability [35, 36, 37, 38], allows us to find the exact expressions of the transmission coefficient, with the boundary conditions and the impurity potentials determining the quantum transport.
Figure 1: Schematic view of a mesoscopic ring with impurities coupled to the one-dimensional ribbon. (a) Illustration of the setup. (b) Details of the quantum ring of radius $a$ with $N$ point-like impurities (dark circles), separated by the angle $\Delta\phi=2\pi/N$, with the potential given by Eq. (1). The coupling angle $\theta$ is defined as the angle between the first impurity on the left with respect to the line connecting the junction and the center of the ring. Plane waves are incident from the left-hand side of the ribbon and are scattered by the ring.
In this work, we investigate the quantum transport and BICs formation in a modified geometry of a ring with a single contact attached to a narrow ribbon (to simplify the coupling effects) in conjunction with the effects of a large number of impurities (Fig. 1). The obtained form of the transmission probability is explicitly shown in Eq. (2). Based on this result, we then demonstrate that in the case of a single impurity, the system can support the BICs that emerge from the Fano profile collapse in the conductance when the impurity position relative to the junction varies (Fig. 3), with the conductance minima explicitly shown in Fig. 2. The mechanism of this effect lies in the destructive interference of the propagating states along the ring and their ensuing scattering from the single impurity. Furthermore, by considering the case of three impurities, we show that the BIC then also emerges, as explicitly displayed in terms of the Fano profile collapse and the corresponding peak in the local density of states (Fig. 4). We finally discuss the general features and the possible relevance of this setup for the manipulation of BICs in electronic transport.
Model and Method. We here consider a quasi-one-dimensional system, a narrow ribbon, with electrons of effective mass $m^{*}$ and momentum $k$, propagating with the usual quadratic dispersion $E(k)=\hbar^{2}k^{2}/(2m^{*})$, in the regime of ballistic transport. The system consists of a single channel coupled to a ring of radius $a$ with $N$ symmetrically distributed identical point-like impurities, with the local potential for the $l^{\rm th}$ impurity of the form
$V^{(l)}(x)=V_{0}\,\delta(x-a\phi_{l}),$ (1)
[…]
We employ the Landauer formalism to address the quantum transport in this setup, as detailed in Sec. S1 of the Supplementary Material (SM). We find an exact expression for the transmission probability ($T$), for any number of the impurities and coupling angle (Sec. S1 of the SM), which reads as
$T=\frac{\left\{U_{N-1}[f(ka)]g(ka,\theta)\right\}^{2}}{\{1-T_{N}[f(ka)]\}^{2}+\left\{U_{N-1}[f(ka)]g(ka,\theta)\right\}^{2}},$ (2)
and yields the corresponding conductance ($G$), as given by the Landauer formula $G=(2e^{2}/h)T$. Here, $T_{N}$ and $U_{N}$ are the order-$N$ Chebyshev polynomials of the first and second kind, respectively, while the functions $f$ and $g$ are defined as
$f(ka)$ $={\rm Re}\left\{\left(1-i\frac{m^{*}V_{0}}{k\hbar^{2}}\right)e^{ika\Delta\phi}\right\}$ (3)
$g(ka,\theta)$ $={\rm Im}\left\{\left(1-i\frac{m^{*}V_{0}}{k\hbar^{2}}\right)e^{ika\Delta\phi}\right\}$
$+\frac{m^{*}V_{0}}{k\hbar^{2}}\cos{[ka(\Delta\phi-2\theta)]}.$ (4)
We now analyze the quantum transport in the case of a single impurity.
Case of a single impurity. The transmission probability [Eq. (2)] for a single impurity takes the form
$T$ =$=\frac{\left\{\sin{[2\pi ka]}-\frac{m^{*}V_{0}}{k\hbar^{2}}\cos{[2\pi ka]}+\frac{m^{*}V_{0}}{k\hbar^{2}}\cos{[ka\left(2\pi-2\theta\right)]}\right\}^{2}}{\left\{1-\cos{[2\pi ka]}-\frac{m^{*}V_{0}}{k\hbar^{2}}\sin{[2\pi ka]}\right\}^{2}+\left\{\sin{[2\pi ka]}-\frac{m^{*}V_{0}}{k\hbar^{2}}\cos{[2\pi ka]}+\frac{m^{*}V_{0}}{k\hbar^{2}}\cos[ka\left(2\pi-2\theta\right)\right\}^{2}}.$ (5)
In Fig. 2, we show the corresponding conductance as a function of the (dimensionless) Fermi wavenumber $ka$ for different local potentials ($V_{0}$) and junction angle ($\theta$). As observed, the conductance oscillates between zero and the conductance quantum $2e^{2}/h$. Furthermore, the resonant states occur at the same (discrete) wavenumbers of the isolated ring with the impurity, and resonances are gradually splitting when the potential strength increases, as found at integer values of $ka$. The formation of asymmetric Fano profiles is observed between the two split resonances, with their characteristic width dependent on the coupling angle ($\theta$).
Fano profiles are fundamentally related to the BICs, although, in practice, they are partially confined leaky modes (quasi-BICs). To get better insight into this phenomenon, we plot the conductance versus the dimensionless wavevector $ka$ and the coupling angle $\theta$ in Fig. 3. Most importantly, it shows that the Fano profile can collapse for certain parameter values in states that can be effectively considered simultaneously resonant and antiresonant. Indeed, at wavenumbers of the isolated ring, the states form standing waves that, in turn, destructively interfere via the coupling (junction) point, acting as a source of confinement, with the conductance taking an indeterminate form. The pattern of even oscillations between resonances adjacent to integer values of $ka$ indicates that, at least for the inversion-symmetric configuration ($\theta=\pi$), all Fano profiles collapse over the entire range of Fermi wavenumbers, as implied by Eq. (5).
To further support our findings, in Fig. S1 of the SM, we display the conductance and local density of states (LDOS) versus the wavenumber when the system approaches the inversion-symmetric configuration, where we explicitly show that at points where the conductance displays the Fano profiles, the LDOS exhibits a narrow profile that reduces its characteristic width when $\theta\to\pi$. This behavior implies that one of the ring states decouples from the delocalized incident ribbon waves while staying localized, generating a BIC in the limit. In the following, we analyze how these results generalize to an arbitrary number of impurities.
Figure 2: Dependence of the conductance $G$ on the Fermi wavenumber $ka$ for a tangential ring with one impurity for different values of the coupling angle $\theta$ and the impurity potential $V_{0}$, as given by Eq. (5). (a) $\theta=0.7\pi$, (b) $\theta=0.8\pi$ and (c) $\theta=0.9\pi$. For a fixed $\theta$, the conductance is shown as the impurity potential, measured in units $\hbar^{2}/m^{*}a$, increases: $V_{0}=0.1$ (black solid line), $V_{0}=0.5$ (red dashed-dotted line), and $V_{0}=0.8$ (blue dashed line). As the potential strength $V_{0}$ is increased, the resonant wavenumbers gradually modify and eventually split, leaving those with integer values of $ka$ unaltered. On the other hand, the change in the coupling angle affects the wavenumbers of anti-resonances and the asymmetric Fano profiles.
Generalization to $N$ impurities. To appreciate the situation with several impurities, we first notice that in a clean, impurity-free ring, the transmission reduces to the well-known expression $T=\cos^{2}(\pi ka)$, which implies that the maxima occur for integer values of $ka$, i.e., for the allowed wavenumbers of an isolated ring. At these wavenumbers, the incident ribbon waves are perfectly transmitted without any phase change after crossing the junction, with only the amplitudes of the ribbon and ring waves being modulated, because of the boundary conditions and the ensuing periodicity. This feature extends in the presence of impurities, since the boundary conditions at the junction do not involve impurities, with the only exception being possible BICs. In the following, we therefore analyze the corresponding spectrum.
For an isolated ring with $N$ pointlike impurities, the spectrum is obtained by applying the Bloch’s theorem, $\psi(x+a\Delta\phi)=e^{iqa\Delta\phi}\psi(x)$, and is determined by the transcendental equation (Sec. S3 of the SM)
$\cos{(qa\Delta\phi)}=\cos{(ka\Delta\phi)}+\frac{mV_{0}}{k\hbar^{2}}\sin{(ka\Delta\phi)}\equiv f(ka),$ (6)
with $qa\in\mathbb{Z}$ in the first Brillouin zone, as dictated by the discrete translational symmetry of the impurity configuration. Notice that the right-hand side is the same function $f(ka)$ obtained in Eq. (2) for the laterally coupled setup. Furthermore, recalling the momentum-energy relation, $k=\sqrt{2m^{*}E}/\hbar$, Eq. (6) can be interpreted as a dispersion relation for the ring states, which implies that the ring spectrum is the same as in the one-dimensional Kronig-Penney model [39, 40]. In particular, the same behavior of allowed and forbidden (resonant) bands appears in our system, with the wavenumbers (energies) for which $|f(ka)|>1$ in Eq. (6), corresponding to the forbidden bands, as it can be seen by considering the dimensionless coupling $v_{0}\equiv V_{0}/(\hbar^{2}/(m^{*}a))$ as a small perturbation ($v_{0}\ll 1$). In the laterally coupled setup, Eq. (6) implies that incident plane wave states at the wavenumber $ka$ that obey it feature the perfect transmission in Eq. (2), $T=1$, by virtue of the property of Chebyshev polynomials, $T_{N}(\cos\alpha)=\cos(N\alpha)$.
Figure 3: Conductance $G$ in units of $2e^{2}/h$ as a function of the coupling angle $\theta$ and the Fermi wavenumber $ka$ for a ring with one impurity and $V_{0}=0.3$ in units of $\hbar^{2}/m^{*}a$. The white arrow indicates the collapse at the lowest energy. Notice that when the minima and maxima of the conductance coincide, the Fano profile collapses.
To analyze the formation of the BICs, the states localized but embedded in the continuum spectrum, we first recall that they do not propagate away from the ring, and the ribbon waves are thus cancelled out. Notice that in a laterally coupled setup, the absence of the transmission implies that the corresponding wavefunction vanishes at the junction point and along the entire right-hand side of the ribbon ($\psi\sim te^{ikx}$), see also Fig. 1(b). Therefore, the BICs can be formed by fixing the junction point so that destructive interference occurs for the wavefunction in a ring and the incident wave. Recall that this effect is possible because the amplitudes of waves in the ring and ribbon are only related by the continuity condition. Then, gradual vanishing of the incident state when approaching such a node ($\sim\theta$), thereby forming a BIC, implies a divergence in the LDOS. This analysis applies as long as the state of the ring is non-degenerate, and therefore pertains to the high-symmetry momenta (at the center and edges) in the Brillouin zone.
Figure 4: (a) Conductance ($G$) in units of $2e^{2}/h$ as a function of the coupling angle $\theta$ and the dimensionless Fermi wavenumber $ka$ for a ring with three impurities and $V_{0}=0.3$, in units of $\hbar^{2}/m^{*}a$. (b) Horizontal line-cuts for $\theta=0.5\Delta\phi$ (red line), $\theta=0.45\Delta\phi$ (blue line), $\theta=0.3\Delta\phi$ (yellow line) and $\theta=0.26\Delta\phi$ (purple line), obtained from Eq. (2), with $\Delta\phi=2\pi/3$ as the angle between the nearest-neighbor impurities. (c) The local density of states for the same values of the parameters.
To illustrate these effects, in Fig. 4(a) we plot the conductance $G$ vs $ka$ and $\theta$ for three impurities and a potential strength $V_{0}=0.3$ in units of $\hbar^{2}/m^{*}a$, showing the band structure and the Fano profile collapse for the momenta $qa=3m$, $m=1,2,...$, at the center of the $(m+1)^{\rm th}$ Brillouin zone. The horizontal cuts for different coupling angles, indicated by the colored lines in Fig. 4(a), are displayed in Fig. 4(b) around the wavevectors corresponding to the collapse of the Fano profile in the conductance. For the destructive interference angles, the collapse of the Fano profile is seen as a disappearance of the resonance, featuring the so-called Ghost Fano effect [9]. On the other hand, Fig. 4(c) shows the LDOS for the same parameters, see also Sec. S2 of the SM for additional details. The divergence and collapse of the delta-like profile of the LDOS verify the decoupling of the state at the corresponding coupling angles.
Further analysis in terms of Bloch functions, indicates that in the limit $V\rightarrow 0$ (isolated impurity-free ring), the splitting of the bands occurs for the symmetric and antisymmetric combinations of clockwise and counterclockwise propagating waves. Moreover, since the number of impurities dictates which states break the degeneracy from $ka\in\mathbb{N}$, eventually forming the BICs, the scatterers can manipulate the electronic transport by decoupling one of the states at the band-edge from the ribbon, which, in turn, drastically modifies the conductance through decoupling of the resonance. Nevertheless, any slight modification would lead to the appearance of the Fano profile and radiation outside the system (quasi-BIC). To undo the latter, additional effects in the ring, such as the Aharonov-Bohm phenomenon and the Rashba spin-orbit interaction, may be employed, but are outside the scope of this work.
Let us now analyze the conductance in the neighborhood of $\tilde{x}_{m}\equiv k_{m}a=mN$, with $m\in\mathbb{N}$, with respect to the potential strength and coupling orientation, near the inversion-symmetric impurity configuration where a BIC can be formed. As a result, the Fano profile is obtained by expanding the transmission coefficient in a series. Defining $\theta=\Delta\phi/2-\Delta/2$, $\tilde{x}=ka$, $v_{0}=m^{*}V_{0}a/\hbar^{2}$ and keeping terms such that ${\rm max}(|\tilde{x}-\tilde{x}_{m}|)\sim\Delta^{2}$, the corresponding transmission probability takes the form
$T\approx\left(\frac{1}{1+q_{F}^{2}}\right)\frac{\left(\tilde{X}-q_{F}\Gamma(\Delta)\right)^{2}}{\tilde{X}^{2}+\Gamma(\Delta)^{2}},$ (7)
where $\tilde{X}(\Delta)\equiv\tilde{x}-\tilde{x}_{m}-\frac{q_{F}(\tilde{x}_{m}\Delta)^{2}}{4\pi}$. Here, the Fano parameter and characteristic width are, respectively, given by
$q_{F}=N\frac{v_{0}}{\tilde{x}_{m}}$    and $\Gamma(\Delta)=\frac{(\tilde{x}_{m}\Delta)^{2}}{4\pi(1+q_{F}^{-2})}.$ (8)
The fact that $\Gamma\sim\Delta^{2}$ indicates that for any number of impurities ($N$) the Fano profile collapses when $\Delta\rightarrow 0$, and implies that the inversion-symmetric configuration ($\theta=\Delta\phi/2$) can host BICs. Therefore, the dependence of the form of the Fano profile on the parameters of the system shows the crucial role of the interference effects in the quantum transport, while the inclusion of the impurities allows for its manipulation.
Conclusions and outlook. We have studied the quantum transport and formation of the BICs in a ring with periodically distributed impurities. Particularly, using Landauer’s formalism, we have derived an exact expression for the conductance [Eq. (2)]. In the case of the tangential coupling between the ring and the ribbon, we emphasize that there is a direct correspondence between the coupled system’s spectrum and the resonant energies [Eq. (6)]. Most importantly, we show that the BICs are formed due to physical conditions that preclude their projection onto the ribbon, therefore implying their decoupling from the continuum spectrum and the concomitant localization. In the case of periodically distributed impurities, the manipulation of the quantum transport is particularly prominent when the inversion symmetry is present, allowing for a selective decoupling of resonant states. Further manipulation may be achieved via the Aharonov-Bohm effect and the Rashba spin-orbit coupling, which we plan to study in future. Finally, the proposed setup could be utilized as an efficient impurity detector due to its extreme sensitivity to the symmetry.
Acknowledgment. This work is supported by the Swedish Research Council Grant No. VR 2019-04735 (V.J.), Fondecyt (Chile) Grants No. 1201876 (P.A.O.) and No. 1230933 (V.J.).

### 27 Feb 2024 v2

@@
 Quantum rings coupled to external nanowires offer a versatile platform for the manipulation of the quantum mesoscopic transport.
-We here study such a system including periodically distributed pointlike impurities along the ring.
-Based on an exact expression for the conductance found here, we demonstrate that the bound-states in continuum (BICs) form from the ring states at the high-symmetry momenta in the ring’s Brillouin zone.
-Furthermore, the presence of the inversion symmetry allows for a selective decoupling of resonant states, favoring the BIC generation and therefore allowing an extra tunability in the quantum transport of the system.
+Here, we study such a system, including periodically distributed pointlike impurities along the ring.
+Based on an exact expression for the conductance found here, we demonstrate that the bound states in the continuum (BICs) form from the ring states at the high-symmetry momenta in the ring’s Brillouin zone.
+Furthermore, the presence of the inversion symmetry allows for a selective decoupling of resonant states, favoring the BIC generation and, therefore, allowing extra tunability in the quantum transport of the system.
 Finally, we suggest that the magnetic fluxes and Rashba spin-orbit coupling offer other possible routes for the BIC formation in laterally coupled quantum rings.
@@
 Bound states in the continuum (BICs) remain localized and coexist with a continuous spectrum of radiating states that propagate outside the system [1, 2].
-Although this special class of states was proposed at the beginning of quantum mechanics by von Neumann and Wigner [3], only in the following decades such states became the subject of more intense research [4, 5].
-Furthermore, the concept of BICs is not exclusively operative in quantum mechanics but also pertains to classical wave phenomena, as they have been shown to occur in various setups, e.g., in photonics, acoustics, and electronics [6, 7, 8, 9, 10].
-Due to the advancement in nanofabrication technology and their experimental observation [11, 12], BICs have recently attracted considerable interest for applications, as in building lasers, sensors, and filters [13, 14, 15, 16, 17, 18].
+Although this special class of states was proposed at the dawn of quantum mechanics by von Neumann and Wigner [3], only in the following decades such states became the subject of more intense research [4, 5].
+Furthermore, the concept of BICs is not exclusively operative in quantum mechanics but also pertains to classical wave phenomena, e.g., in photonic, acoustic, and electronic setups [6, 7, 8, 9, 10, 11, 12, 13, 14].
+Due to advancements in nanofabrication technology and concomitant experimental observation of the BIC and BIC-like states in these systems [15, 16, 17, 18, 19, 20, 21], they have also recently attracted considerable interest for applications, as in building lasers, photonic circuit elements, sensors, and filters [22, 23, 24, 25, 26, 27, 28].
 In mesoscopic systems, BICs can be formed by several mechanisms, with destructive interference in resonant states possibly being the most prominent one [1].
-In this respect, quantum rings play a rather special role due to their geometry, also offering flexibility of manipulation, as they have already been intensively studied both theoretically [19, 20, 21, 22, 23, 24, 25] and experimentally [26, 27, 28, 29, 30, 31, 32, 33].
+In this respect, quantum rings play a rather special role due to their geometry, also offering flexibility of manipulation, as they have already been intensively studied both theoretically [29, 30, 31, 32, 33, 34, 35, 36, 37] and experimentally [38, 39, 40, 41, 42, 43, 44].
 Moreover, mesoscopic coupled waveguide systems feature various sources of disorder, often modeled as point-like impurities or short-range scatterers [29, 31, 32, 45].
 On the other hand, in such coupled mesoscopic systems, the appearance of a subset of discrete energy levels mixed with a continuum of states is attributed to the Fano effect [46], manifesting as an asymmetric peak-and-dip feature in the conductance.
-However, the case of many impurities has not been intensively studied despite the expected richness in the transport behavior.
-In particular, a strong dependence of the conductance on the distribution of impurities and the system’s geometry is expected primarily due to quantum interference phenomena dominating the physics at the mesoscopic scales.
-Furthermore, the theory of electronic transport, relating the conductance to the transmission probability [35, 36, 37, 38], allows us to find the exact expressions of the transmission coefficient, with the boundary conditions and the impurity potentials determining the quantum transport.
+However, the case of many impurities has not been intensively explored in spite of the expected richness in the transport behavior.
+In particular, the behavior of the conductance should be strongly dependent on the impurity distribution and system’s geometry, owing to quantum interference effects at mesoscopic scales.
+Furthermore, the theory of electronic transport, relating the conductance to the transmission probability [47, 48, 49, 50], allows us to find the exact expressions of the transmission coefficient, with the boundary conditions and the impurity potentials determining the quantum transport therein.
 Figure 1: Schematic view of a mesoscopic ring with impurities coupled to the one-dimensional ribbon.
@@
 To further support our findings, in Fig.
-S1 of the SM, we display the conductance and local density of states (LDOS) versus the wavenumber when the system approaches the inversion-symmetric configuration, where we explicitly show that at points where the conductance displays the Fano profiles, the LDOS exhibits a narrow profile that reduces its characteristic width when $\theta\to\pi$.
+S1 of the SM, we display the conductance and local density of states (LDOS) versus the wavenumber when the system approaches the inversion-symmetric configuration, where we explicitly show that at points where the conductance displays the Fano profiles, the local density of states exhibits a narrow profile that reduces its characteristic width when $\theta\to\pi$.
 This behavior implies that one of the ring states decouples from the delocalized incident ribbon waves while staying localized, generating a BIC in the limit.
@@
 To appreciate the situation with several impurities, we first notice that in a clean, impurity-free ring, the transmission reduces to the well-known expression $T=\cos^{2}(\pi ka)$, which implies that the maxima occur for integer values of $ka$, i.e., for the allowed wavenumbers of an isolated ring.
-At these wavenumbers, the incident ribbon waves are perfectly transmitted without any phase change after crossing the junction, with only the amplitudes of the ribbon and ring waves being modulated, because of the boundary conditions and the ensuing periodicity.
-This feature extends in the presence of impurities, since the boundary conditions at the junction do not involve impurities, with the only exception being possible BICs.
+At these wavenumbers, the incident ribbon waves are perfectly transmitted without any phase change after crossing the junction, with the amplitudes of the ribbon and ring waves being modulated because of the boundary conditions and the ensuing periodicity.
+This feature extends in the presence of impurities since the boundary conditions at the junction do not involve impurities, with the only exception being possible BICs.
 In the following, we therefore analyze the corresponding spectrum.
@@
 In the laterally coupled setup, Eq.
-(6) implies that incident plane wave states at the wavenumber $ka$ that obey it feature the perfect transmission in Eq.
+(6) implies that incident plane wave states (with the wavenumber $ka$) obeying it, feature the perfect transmission in Eq.
 (2), $T=1$, by virtue of the property of Chebyshev polynomials, $T_{N}(\cos\alpha)=\cos(N\alpha)$.
@@
 Notice that when the minima and maxima of the conductance coincide, the Fano profile collapses.
-To analyze the formation of the BICs, the states localized but embedded in the continuum spectrum, we first recall that they do not propagate away from the ring, and the ribbon waves are thus cancelled out.
+To analyze the formation of the BICs, we first recall that they do not propagate away from the ring, and the ribbon waves are thus canceled out.
 Notice that in a laterally coupled setup, the absence of the transmission implies that the corresponding wavefunction vanishes at the junction point and along the entire right-hand side of the ribbon ($\psi\sim te^{ikx}$), see also Fig.
@@
 Recall that this effect is possible because the amplitudes of waves in the ring and ribbon are only related by the continuity condition.
-Then, gradual vanishing of the incident state when approaching such a node ($\sim\theta$), thereby forming a BIC, implies a divergence in the LDOS.
+Then, the gradual vanishing of the incident state when approaching such a node ($\sim\theta$), thereby forming a BIC, implies a divergence in the LDOS.
 This analysis applies as long as the state of the ring is non-degenerate, and therefore pertains to the high-symmetry momenta (at the center and edges) in the Brillouin zone.
@@
 On the other hand, Fig.
-4(c) shows the LDOS for the same parameters, see also Sec.
+4(c) shows the LDOS for the same parameters; see also Sec.
 S2 of the SM for additional details.
 The divergence and collapse of the delta-like profile of the LDOS verify the decoupling of the state at the corresponding coupling angles.
-Further analysis in terms of Bloch functions, indicates that in the limit $V\rightarrow 0$ (isolated impurity-free ring), the splitting of the bands occurs for the symmetric and antisymmetric combinations of clockwise and counterclockwise propagating waves.
+Further analysis in terms of Bloch functions indicates that in the limit $V\rightarrow 0$ (isolated impurity-free ring), the splitting of the bands occurs for the symmetric and antisymmetric combinations of clockwise and counterclockwise propagating waves.
 Moreover, since the number of impurities dictates which states break the degeneracy from $ka\in\mathbb{N}$, eventually forming the BICs, the scatterers can manipulate the electronic transport by decoupling one of the states at the band-edge from the ribbon, which, in turn, drastically modifies the conductance through decoupling of the resonance.
 Nevertheless, any slight modification would lead to the appearance of the Fano profile and radiation outside the system (quasi-BIC).
-To undo the latter, additional effects in the ring, such as the Aharonov-Bohm phenomenon and the Rashba spin-orbit interaction, may be employed, but are outside the scope of this work.
+To undo the latter, effects yielding additional phase factors for the wavefunctions on the ring, such as the magnetic Aharonov-Bohm fluxes and the Rashba spin-orbit interaction, may be employed but are outside the scope of this work.
 Let us now analyze the conductance in the neighborhood of $\tilde{x}_{m}\equiv k_{m}a=mN$, with $m\in\mathbb{N}$, with respect to the potential strength and coupling orientation, near the inversion-symmetric impurity configuration where a BIC can be formed.
@@
 $T\approx\left(\frac{1}{1+q_{F}^{2}}\right)\frac{\left(\tilde{X}-q_{F}\Gamma(\Delta)\right)^{2}}{\tilde{X}^{2}+\Gamma(\Delta)^{2}},$ (7)
-where $\tilde{X}(\Delta)\equiv\tilde{x}-\tilde{x}_{m}-\frac{q_{F}(\tilde{x}_{m}\Delta)^{2}}{4\pi}$. Here, the Fano parameter and characteristic width are, respectively, given by
+where $\tilde{X}(\Delta)\equiv\tilde{x}-\tilde{x}_{m}-\frac{q_{F}(\tilde{x}_{m}\Delta)^{2}}{4\pi(1+q_{F}^{2})}$. Here, the Fano parameter and characteristic width are, respectively, given by
 $q_{F}=N\frac{v_{0}}{\tilde{x}_{m}}$    and $\Gamma(\Delta)=\frac{(\tilde{x}_{m}\Delta)^{2}}{4\pi(1+q_{F}^{-2})}.$ (8)
-The fact that $\Gamma\sim\Delta^{2}$ indicates that for any number of impurities ($N$) the Fano profile collapses when $\Delta\rightarrow 0$, and implies that the inversion-symmetric configuration ($\theta=\Delta\phi/2$) can host BICs.
+The fact that $\Gamma\sim\Delta^{2}$ indicates that for any number of impurities ($N$), the Fano profile collapses when $\Delta\rightarrow 0$, and implies that the inversion-symmetric configuration ($\theta=\Delta\phi/2$) can host BICs.
 Therefore, the dependence of the form of the Fano profile on the parameters of the system shows the crucial role of the interference effects in the quantum transport, while the inclusion of the impurities allows for its manipulation.
@@
 In the case of periodically distributed impurities, the manipulation of the quantum transport is particularly prominent when the inversion symmetry is present, allowing for a selective decoupling of resonant states.
-Further manipulation may be achieved via the Aharonov-Bohm effect and the Rashba spin-orbit coupling, which we plan to study in future.
-Finally, the proposed setup could be utilized as an efficient impurity detector due to its extreme sensitivity to the symmetry.
+Further manipulation may be achieved via the Aharonov-Bohm effect and the Rashba spin-orbit coupling, which we plan to study in the future.
+Finally, the proposed setup could be utilized as an efficient impurity detector due to its extreme sensitivity to symmetry.
 Acknowledgment. This work is supported by the Swedish Research Council Grant No. VR 2019-04735 (V.J.), Fondecyt (Chile) Grants No. 1201876 (P.A.O.) and No. 1230933 (V.J.).